# TopoBench: A Framework for Benchmarking Topological Deep Learning

### v2

@@ C.4 Additional Results
Dataset CCXN CWN CCCN SCCN SCCNN SCN DR SDP DR SDP DR SDP DR SDP DR SDP DR SDP Node-level tasks Cora 86.32 ± 1.22 86.79 ± 1.81 74.95 ± 0.98 86.32 ± 1.38 87.44 ± 1.28 87.68 ± 1.17 80.86 ± 2.16 80.06 ± 1.66 82.19 ± 1.07 80.65 ± 2.39 82.27 ± 1.34 79.91 ± 1.18 Citeseer 72.87 ± 1.13 74.67 ± 2.24 70.49 ± 2.85 75.2 ± 1.82 75.63 ± 1.58 74.91 ± 1.25 69.6 ± 1.83 68.86 ± 2.40 70.23 ± 2.69 69.03 ± 2.01 71.24 ± 1.68 70.4 ± 1.53 Pubmed 88.91 ± 0.47 88.38 ± 0.38 86.94 ± 0.68 88.64 ± 0.36 88.52 ± 0.44 88.67 ± 0.39 88.04 ± 0.51 88.37 ± 0.48 88.18 ± 0.32 87.78 ± 0.58 88.72 ± 0.5 88.62 ± 0.44 Amazon 48.93 ± 0.14 48.34 ± 0.12 45.58 ± 0.25 51.9 ± 0.15 50.55 ± 0.15 50.26 ± 0.17 OOM OOM OOM OOM OOM OOM Empire 80.46 ± 0.23 81.44 ± 0.31 66.13 ± 0.03 81.81 ± 0.62 82.14 ± 0.00 82.51 ± 0.0 88.2 ± 0.22 88.27 ± 0.14 89.15 ± 0.32 88.73 ± 0.12 85.89 ± 0.34 88.79 ± 0.46 Minesweeper $8$ 0.36 89.76 ± 0.32 48.82 ± 0.0 88.62 ± 0.04 89.42 ± 0.00 89.85 ± 0.00 88.85 ± 0.00 89.07 ± 0.25 87.4 ± 0.0 89.0 ± 0.00 90.32 ± 0.11 90.27 ± 0.36 Election 0.39 ± 0.05 0.35 ± 0.02 0.6 ± 0.04 0.34 ± 0.02 0.31 ± 0.02 0.31 ± 0.01 0.53 ± 0.03 0.57 ± 0.02 0.51 ± 0.03 0.56 ± 0.04 0.46 ± 0.04 0.51 ± 0.03 Bachelor 0.33 ± 0.03 0.32 ± 0.03 0.33 ± 0.03 0.33 ± 0.03 0.32 ± 0.02 0.31 ± 0.02 0.36 ± 0.02 0.34 ± 0.02 0.34 ± 0.03 0.34 ± 0.03 0.32 ± 0.02 0.32 ± 0.03 Birth 0.80 ± 0.12 0.74 ± 0.11 0.81 ± 0.11 0.72 ± 0.09 0.71 ± 0.09 0.72 ± 0.05 0.82 ± 0.09 0.83 ± 0.10 0.79 ± 0.12 0.83 ± 0.12 0.71 ± 0.08 0.8 ± 0.11 Death 0.57 ± 0.06 0.54 ± 0.06 0.55 ± 0.05 0.54 ± 0.06 0.54 ± 0.06 0.54 ± 0.06 0.58 ± 0.06 0.56 ± 0.04 0.55 ± 0.05 0.58 ± 0.05 0.52 ± 0.05 0.56 ± 0.05 Income 0.25 ± 0.03 0.25 ± 0.03 0.36 ± 0.04 0.25 ± 0.03 0.23 ± 0.02 0.23 ± 0.02 0.29 ± 0.03 0.29 ± 0.03 0.28 ± 0.03 0.31 ± 0.03 0.25 ± 0.02 0.27 ± 0.02 Migration 0.80 ± 0.11 0.85 ± 0.18 0.9 ± 0.16 0.84 ± 0.13 0.84 ± 0.10 0.84 ± 0.12 0.91 ± 0.18 0.93 ± 0.17 0.90 ± 0.14 0.93 ± 0.17 0.92 ± 0.20 0.96 ± 0.23 Unempl 0.28 ± 0.05 0.27 ± 0.03 0.46 ± 0.04 0.25 ± 0.03 0.24 ± 0.03 0.25 ± 0.03 0.43 ± 0.04 0.47 ± 0.04 0.43 ± 0.04 0.45 ± 0.04 0.38 ± 0.04 0.41 ± 0.03 Graph-level tasks MUTAG 69.79 ± 4.61 74.89 ± 5.51 69.68 ± 8.58 80.43 ± 1.78 80.85 ± 5.42 77.02 ± 9.32 70.64 ± 5.90 73.62 ± 4.41 76.17 ± 6.63 70.64 ± 3.16 71.49 ± 2.43 73.62 ± 6.13 PROTEINS 75.63 ± 2.57 74.91 ± 1.85 76.13 ± 1.80 76.13 ± 2.70 73.55 ± 3.43 73.33 ± 2.30 75.05 ± 2.76 74.34 ± 3.17 74.19 ± 2.86 74.98 ± 1.92 75.27 ± 2.14 74.77 ± 1.69 NCI1 72.43 ± 1.72 74.86 ± 0.82 68.52 ± 0.51 73.93 ± 1.87 76.67 ± 1.48 77.65 ± 1.28 76.42 ± 0.88 76.17 ± 1.39 76.6 ± 1.75 75.6 ± 2.45 75.27 ± 1.57 74.49 ± 1.03 NCI109 73.22 ± 0.48 75.66 ± 1.30 68.19 ± 0.65 73.8 ± 2.06 75.35 ± 1.50 74.83 ± 1.18 75.49 ± 1.39 75.31 ± 1.36 77.12 ± 1.07 75.43 ± 1.94 74.58 ± 1.29 75.7 ± 1.04 IMDB-BIN 70.08 ± 1.21 68.96 ± 2.03 70.4 ± 2.02 69.28 ± 2.57 69.12 ± 2.82 69.44 ± 2.46 70.88 ± 3.98 69.76 ± 3.16 70.88 ± 2.25 69.28 ± 5.69 70.8 ± 2.38 68.64 ± 3.90 IMDB-MUL 47.63 ± 3.45 48.75 ± 3.56 49.71 ± 2.83 49.87 ± 2.33 49.01 ± 2.63 47.79 ± 3.45 49.71 ± 3.70 47.31 ± 3.12 48.75 ± 3.98 46.67 ± 3.13 48.16 ± 2.89 49.49 ± 5.08 REDDIT 74.40 ± 1.50 82.84 ± 2.54 76.20 ± 0.86 85.52 ± 1.38 85.12 ± 1.29 83.32 ± 0.73 74.16 ± 1.54 74.44 ± 1.74 75.56 ± 3.46 77.24 ± 1.87 71.28 ± 2.06 69.68 ± 4.0 ZINC 0.63 ± 0.02 0.40 ± 0.04 0.70 ± 0.0 0.34 ± 0.01 0.35 ± 0.02 0.34 ± 0.02 0.55 ± 0.01 0.46 ± 0.08 0.36 ± 0.01 0.36 ± 0.02 0.59 ± 0.01 0.53 ± 0.04

### v3

@@ C.4 Additional Results and Analysis
Dataset CCXN CWN CCCN SCCN SCCNN SCN DR SDP DR SDP DR SDP DR SDP DR SDP DR SDP Node-level tasks Cora 86.32 ± 1.22 86.79 ± 1.81 74.95 ± 0.98 86.32 ± 1.38 87.44 ± 1.28 87.68 ± 1.17 80.86 ± 2.16 80.06 ± 1.66 82.19 ± 1.07 80.65 ± 2.39 82.27 ± 1.34 79.91 ± 1.18 Citeseer 72.87 ± 1.13 74.67 ± 2.24 70.49 ± 2.85 75.2 ± 1.82 75.63 ± 1.58 74.91 ± 1.25 69.6 ± 1.83 68.86 ± 2.40 70.23 ± 2.69 69.03 ± 2.01 71.24 ± 1.68 70.4 ± 1.53 Pubmed 88.91 ± 0.47 88.38 ± 0.38 86.94 ± 0.68 88.64 ± 0.36 88.52 ± 0.44 88.67 ± 0.39 88.04 ± 0.51 88.37 ± 0.48 88.18 ± 0.32 87.78 ± 0.58 88.72 ± 0.5 88.62 ± 0.44 Amazon 48.93 ± 0.14 48.34 ± 0.12 45.58 ± 0.25 51.9 ± 0.15 50.55 ± 0.15 50.26 ± 0.17 OOM OOM OOM OOM OOM OOM Empire 80.46 ± 0.23 81.44 ± 0.31 66.13 ± 0.03 81.81 ± 0.62 82.14 ± 0.00 82.51 ± 0.0 88.2 ± 0.22 88.27 ± 0.14 89.15 ± 0.32 88.73 ± 0.12 85.89 ± 0.34 88.79 ± 0.46 Minesweeper 88.88 ± 0.36 89.76 ± 0.32 48.82 ± 0.0 88.62 ± 0.04 89.42 ± 0.00 89.85 ± 0.00 88.85 ± 0.00 89.07 ± 0.25 87.4 ± 0.0 89.0 ± 0.00 90.32 ± 0.11 90.27 ± 0.36 Election 0.39 ± 0.05 0.35 ± 0.02 0.6 ± 0.04 0.34 ± 0.02 0.31 ± 0.02 0.31 ± 0.01 0.53 ± 0.03 0.57 ± 0.02 0.51 ± 0.03 0.56 ± 0.04 0.46 ± 0.04 0.51 ± 0.03 Bachelor 0.33 ± 0.03 0.32 ± 0.03 0.33 ± 0.03 0.33 ± 0.03 0.32 ± 0.02 0.31 ± 0.02 0.36 ± 0.02 0.34 ± 0.02 0.34 ± 0.03 0.34 ± 0.03 0.32 ± 0.02 0.32 ± 0.03 Birth 0.80 ± 0.12 0.74 ± 0.11 0.81 ± 0.11 0.72 ± 0.09 0.71 ± 0.09 0.72 ± 0.05 0.82 ± 0.09 0.83 ± 0.10 0.79 ± 0.12 0.83 ± 0.12 0.71 ± 0.08 0.8 ± 0.11 Death 0.57 ± 0.06 0.54 ± 0.06 0.55 ± 0.05 0.54 ± 0.06 0.54 ± 0.06 0.54 ± 0.06 0.58 ± 0.06 0.56 ± 0.04 0.55 ± 0.05 0.58 ± 0.05 0.52 ± 0.05 0.56 ± 0.05 Income 0.25 ± 0.03 0.25 ± 0.03 0.36 ± 0.04 0.25 ± 0.03 0.23 ± 0.02 0.23 ± 0.02 0.29 ± 0.03 0.29 ± 0.03 0.28 ± 0.03 0.31 ± 0.03 0.25 ± 0.02 0.27 ± 0.02 Migration 0.80 ± 0.11 0.85 ± 0.18 0.9 ± 0.16 0.84 ± 0.13 0.84 ± 0.10 0.84 ± 0.12 0.91 ± 0.18 0.93 ± 0.17 0.90 ± 0.14 0.93 ± 0.17 0.92 ± 0.20 0.96 ± 0.23 Unempl 0.28 ± 0.05 0.27 ± 0.03 0.46 ± 0.04 0.25 ± 0.03 0.24 ± 0.03 0.25 ± 0.03 0.43 ± 0.04 0.47 ± 0.04 0.43 ± 0.04 0.45 ± 0.04 0.38 ± 0.04 0.41 ± 0.03 Graph-level tasks MUTAG 69.79 ± 4.61 74.89 ± 5.51 69.68 ± 8.58 80.43 ± 1.78 80.85 ± 5.42 77.02 ± 9.32 70.64 ± 5.90 73.62 ± 4.41 76.17 ± 6.63 70.64 ± 3.16 71.49 ± 2.43 73.62 ± 6.13 PROTEINS 75.63 ± 2.57 74.91 ± 1.85 76.13 ± 1.80 76.13 ± 2.70 73.55 ± 3.43 73.33 ± 2.30 75.05 ± 2.76 74.34 ± 3.17 74.19 ± 2.86 74.98 ± 1.92 75.27 ± 2.14 74.77 ± 1.69 NCI1 72.43 ± 1.72 74.86 ± 0.82 68.52 ± 0.51 73.93 ± 1.87 76.67 ± 1.48 77.65 ± 1.28 76.42 ± 0.88 76.17 ± 1.39 76.6 ± 1.75 75.6 ± 2.45 75.27 ± 1.57 74.49 ± 1.03 NCI109 73.22 ± 0.48 75.66 ± 1.30 68.19 ± 0.65 73.8 ± 2.06 75.35 ± 1.50 74.83 ± 1.18 75.49 ± 1.39 75.31 ± 1.36 77.12 ± 1.07 75.43 ± 1.94 74.58 ± 1.29 75.7 ± 1.04 IMDB-BIN 70.08 ± 1.21 68.96 ± 2.03 70.4 ± 2.02 69.28 ± 2.57 69.12 ± 2.82 69.44 ± 2.46 70.88 ± 3.98 69.76 ± 3.16 70.88 ± 2.25 69.28 ± 5.69 70.8 ± 2.38 68.64 ± 3.90 IMDB-MUL 47.63 ± 3.45 48.75 ± 3.56 49.71 ± 2.83 49.87 ± 2.33 49.01 ± 2.63 47.79 ± 3.45 49.71 ± 3.70 47.31 ± 3.12 48.75 ± 3.98 46.67 ± 3.13 48.16 ± 2.89 49.49 ± 5.08 REDDIT 74.40 ± 1.50 82.84 ± 2.54 76.20 ± 0.86 85.52 ± 1.38 85.12 ± 1.29 83.32 ± 0.73 74.16 ± 1.54 74.44 ± 1.74 75.56 ± 3.46 77.24 ± 1.87 71.28 ± 2.06 69.68 ± 4.0 ZINC 0.63 ± 0.02 0.40 ± 0.04 0.70 ± 0.0 0.34 ± 0.01 0.35 ± 0.02 0.34 ± 0.02 0.55 ± 0.01 0.46 ± 0.08 0.36 ± 0.01 0.36 ± 0.02 0.59 ± 0.01 0.53 ± 0.04
@@
Table 9: Model runtime in seconds corresponding to the best set of hyperparameters
Model GCN GAT GIN AST EDGNN UniGNN2 CWN CCCN CCXN SCN SCCN SCCNN Cora 19.71 ± 3.18 33.51 ± 6.33 24.24 ± 7.48 40.23 ± 11.82 28.17 ± 3.97 35.28 ± 5.14 48.49 ± 27.58 39.36 ± 6.72 35.74 ± 2.93 51.53 ± 14.0 84.04 ± 13.39 65.62 ± 23.49 Citeseer 22.58 ± 2.48 21.18 ± 2.24 24.65 ± 2.58 58.78 ± 2.26 39.59 ± 1.93 41.58 ± 3.55 48.89 ± 4.02 48.39 ± 4.57 53.88 ± 2.81 58.85 ± 18.23 65.53 ± 13.34 80.6 ± 36.33 Pubmed 40.89 ± 8.39 50.2 ± 13.27 59.91 ± 10.56 84.67 ± 13.03 87.25 ± 13.88 160.37 ± 31.88 147.59 ± 28.96 172.63 ± 33.06 212.66 ± 52.26 151.0 ± 22.55 134.82 ± 34.81 171.92 ± 48.25 MUTAG 3.83 ± 0.89 4.16 ± 1.05 4.6 ± 0.56 9.9 ± 3.24 5.81 ± 1.13 5.5 ± 0.78 10.92 ± 0.96 12.04 ± 2.21 10.76 ± 1.89 8.47 ± 2.43 10.71 ± 2.92 14.06 ± 2.51 PROTEINS 8.18 ± 2.47 8.18 ± 2.3 8.88 ± 2.34 15.81 ± 2.89 15.15 ± 3.3 27.8 ± 7.77 53.6 ± 17.7 41.63 ± 7.23 51.98 ± 8.5 42.78 ± 13.41 70.06 ± 15.91 54.13 ± 11.27 NCI1 53.23 ± 19.67 57.32 ± 17.49 61.2 ± 23.97 138.13 ± 46.01 110.86 ± 27.75 169.17 ± 11.25 302.34 ± 63.44 372.36 ± 109.47 244.72 ± 46.09 276.32 ± 63.21 332.76 ± 51.89 307.2 ± 83.01 NCI109 37.4 ± 8.63 56.44 ± 9.05 50.32 ± 7.98 138.66 ± 26.38 126.61 ± 45.53 61.25 ± 16.58 294.79 ± 46.27 272.3 ± 20.89 225.72 ± 86.57 226.23 ± 66.29 321.76 ± 55.92 353.69 ± 105.89 IMDB-BIN 8.18 ± 3.33 7.48 ± 2.21 7.72 ± 1.72 21.34 ± 2.96 13.06 ± 4.76 20.75 ± 6.02 61.16 ± 9.33 51.88 ± 12.3 51.2 ± 18.14 374.83 ± 132.68 432.23 ± 57.29 515.13 ± 112.53 IMDB-MUL 10.27 ± 3.58 9.79 ± 1.85 10.42 ± 3.93 20.89 ± 5.05 13.89 ± 3.33 16.74 ± 4.15 39.99 ± 5.18 73.6 ± 20.07 72.3 ± 13.76 776.65 ± 147.42 716.44 ± 232.31 895.77 ± 399.41 REDDIT 16.17 ± 1.75 26.87 ± 5.02 28.75 ± 7.64 72.33 ± 18.33 73.92 ± 26.58 307.4 ± 159.71 1230.53 ± 270.03 1653.47 ± 641.13 1435.07 ± 427.2 985.1 ± 68.42 1622.17 ± 667.71 3670.64 ± 423.25 Amazon 22.12 ± 4.99 25.09 ± 4.41 15.25 ± 2.73 61.77 ± 8.17 52.31 ± 4.88 264.24 ± 29.28 239.31 ± 45.79 149.34 ± 58.59 201.99 ± 38.94 OOM OOM OOM Minesweeper 7.33 ± 1.66 10.64 ± 1.1 8.32 ± 2.51 15.79 ± 2.67 20.4 ± 4.38 58.72 ± 14.39 52.12 ± 13.87 98.2 ± 15.21 50.27 ± 14.89 82.5 ± 19.14 28.76 ± 3.74 54.1 ± 20.66 Empire 23.38 ± 2.21 27.1 ± 1.89 22.36 ± 4.9 82.61 ± 15.26 63.77 ± 13.16 122.23 ± 35.71 67.26 ± 10.32 84.54 ± 6.34 63.27 ± 15.66 69.98 ± 14.33 94.94 ± 27.26 96.45 ± 40.6 Tolokers 163.93 ± 31.44 119.87 ± 42.98 77.75 ± 26.74 117.83 ± 23.21 210.78 ± 30.25 412.53 ± 94.18 OOM OOM OOM OOM OOM OOM Election 2.27 ± 0.29 2.66 ± 0.43 2.1 ± 0.33 3.33 ± 0.34 2.21 ± 0.34 2.97 ± 0.32 4.01 ± 0.39 4.15 ± 0.28 3.98 ± 0.42 5.18 ± 0.48 4.23 ± 0.73 9.35 ± 4.52 Bachelor 2.28 ± 0.11 2.69 ± 0.39 1.79 ± 0.28 3.09 ± 0.31 2.24 ± 0.38 2.3 ± 0.25 2.99 ± 0.34 3.01 ± 0.32 3.45 ± 0.29 5.11 ± 0.67 5.98 ± 1.8 7.84 ± 2.14 Birth 2.04 ± 0.27 2.29 ± 0.33 1.93 ± 0.32 2.18 ± 0.32 2.2 ± 0.39 2.33 ± 0.29 4.11 ± 0.48 3.03 ± 0.3 4.04 ± 0.45 4.99 ± 0.4 10.3 ± 5.81 29.24 ± 16.97 Death 2.21 ± 0.3 2.1 ± 0.32 1.77 ± 0.32 3.13 ± 0.24 2.35 ± 0.22 2.66 ± 0.46 4.54 ± 0.55 4.08 ± 0.51 3.49 ± 0.32 4.62 ± 0.55 10.31 ± 5.89 8.2 ± 1.82 Income 2.21 ± 0.18 2.66 ± 0.41 2.03 ± 0.35 3.33 ± 0.35 2.6 ± 0.33 2.94 ± 0.37 3.36 ± 0.28 3.71 ± 0.28 3.42 ± 0.31 5.08 ± 0.63 10.51 ± 8.3 5.57 ± 0.7 Migration 1.86 ± 0.3 2.17 ± 0.3 1.87 ± 0.31 3.49 ± 0.24 2.19 ± 0.3 2.69 ± 0.43 4.6 ± 0.35 4.61 ± 0.36 4.48 ± 0.79 5.22 ± 0.47 4.15 ± 0.76 7.79 ± 2.02 Unempl 2.26 ± 0.35 2.68 ± 0.3 2.02 ± 0.31 4.73 ± 0.31 2.17 ± 0.28 4.06 ± 0.08 3.97 ± 0.3 4.19 ± 0.29 5.1 ± 0.8 5.37 ± 0.44 4.17 ± 0.73 5.6 ± 0.67 ZINC 146.89 ± 27.63 171.15 ± 64.67 168.43 ± 107.1 397.77 ± 153.75 357.68 ± 46.95 520.58 ± 235.2 1390.21 ± 96.86 1621.11 ± 141.72 1226.67 ± 317.55 1209.83 ± 571.93 1226.19 ± 1686.57 2060.2 ± 408.2 Average runtime 27.23 ± 43.95 29.37 ± 42.38 26.65 ± 38.89 59.08 ± 88.18 56.14 ± 85.81 102.00 ± 147.61 191.18 ± 384.52 224.64 ± 479.47 188.77 ± 389.49 217.48 ± 355.56 259.50 ± 444.10 420.65 ± 903.88
@@
Appendix D Higher-Order Datasets
D.1 Descriptive Summaries of Higher-Order Datasets
Tables 11 and 12 provide descriptive summaries of the higher-order datasets included in TopoBench, which spans 13 datasets drawn from a broad range of hypergraph and simplicial benchmark sources.
Hypergraph datasets: For the co-authorship networks (Cora-CA and DBLP-CA) and co-citation networks (Cora, Citeseer, and Pubmed), we use the versions provided by Yadati et al. (2019). For 3D object classification, we include the Princeton ModelNet40 (Wu et al., 2015) and National Taiwan University (Chen et al., 2003) datasets, with hypergraphs constructed following the protocols in Feng et al. (2019) and Yang et al. (2020), using both MVCNN (Su et al., 2015) and GVCNN (Feng et al., 2019) features. Additionally, we evaluate performance on three datasets with categorical attributes—20Newsgroups, Mushroom, and ZOO—sourced from the UCI Categorical Machine Learning Repository (Dua et al., 2017). For these, we construct hypergraphs as in Yadati et al. (2019), where a hyperedge is formed by grouping data points sharing the same categorical feature value.
Table 11 summarizes the statistics of the hypergraph datasets along with their associated homophily metrics: clique-expansion homophily (Wang et al., 2023) and $\Delta$-homophily (Telyatnikov et al., 2025). These measures capture the degree to which hyperedges align with label information, serving as indicators of how well the hypergraph structure supports downstream classification.
Simplicial datasets: In contrast, the MANTRA family (Ballester et al., 2024) comprises purely topological datasets of 2-manifold triangulations. From the suite of tasks it offers, we focus on three representative classification problems: (1) NAME: predicting the homeomorphism class of a triangulated surface, (2) ORIENT: determining its orientability, and (3) $\beta_{1}$, $\beta_{2}$: predicting the values of the first and second Betti numbers. The task of predicting Betti numbers is performed as a regression, while the outputs are rounded, and then classification metrics are employed to assess performance. Table 12 reports the corresponding dataset statistics for this family.
Table 11: Statistics of hypergraph higher-order datasets
Dataset 0-cell dim Num. 0-cell Num. $1$-cell Num. classes CE Homophily $\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}h^{0}_{v}$ $\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}h^{1}_{v}$ Cora 1433 2708 1579 7 89.74 84.10 78.08 Citeseer 3703 3312 1079 6 89.32 78.25 74.18 Pubmed 500 19717 7963 3 95.24 82.05 75.73 CORA-CA 1433 2708 1072 7 80.26 80.81 76.51 DBLP-CA 1425 41302 22363 6 86.88 88.86 86.01 ZOO 16 101 43 7 82.88 91.13 85.79 20Newsgroups 100 16262 100 4 75.25 81.26 74.78 Mushroom 22 8124 298 2 85.33 88.05 84.41 NTU2012 100 16242 2012 67 46.07 53.24 41.95 ModelNet40 100 12311 12311 40 24.07 42.16 29.42
Table 12: Statistics of MANTRA family simplicial datasets
Dataset 0/1/2-cell dim Avg. num. 0-cell Avg. num. 1-cell Avg. num. 2-cell Num. classes Num. objects NAME 1/1/1 9.98 34.43 22.95 8 43138 ORIENT 1/1/1 9.98 34.43 22.95 2 43138 $\beta_{1}$$\beta_{2}$ 1/1/1 9.98 34.43 22.95 1 43138
D.2 Hypergraph Higher-Order Datasets Results
Table 13 demonstrates varying model effectiveness across real-world classification hypergraph datasets. While no single model consistently outperforms others across all datasets, AllSetTransformer achieves the best performance in 5 out of 10 cases. UniGNN2 achieves top performance on several datasets, including Cora and ModelNet40, while EDGNN leads on CORA-CA and Citeseer. It is important to note that the results shown in Table 13 and Table 1 for Cora, Citeseer, and Pubmed refer to the same base datasets but differ in the nature of their topology. Specifically, the higher-order structures in Table 1 are derived via lifting mechanisms applied to graph data (graph representation of the Cora, Citeseer, and Pubmed), whereas the results in Table 13 are obtained from real hypergraph datasets, where hyperedges are constructed based on available metadata, please refer to Appendix D.2 of Telyatnikov et al. (2025).
Table 13: Test accuracy (mean $\pm$ std) for each hypergraph dataset (rows) and model (columns). The best result is bold and shaded in gray, while those within one standard deviation are in blue-shaded boxes.
EDGNN AllSetTransformer UniGNN2 Cora 78.14 ± 0.72 78.91 ± 1.06 79.56 ± 1.54 Citeseer 72.58 ± 1.51 71.57 ± 1.71 72.39 ± 2.38 Pubmed 87.04 ± 0.34 87.22 ± 0.28 86.93 ± 0.53 CORA-CA 82.36 ± 0.72 82.19 ± 2.61 81.71 ± 1.42 DBLP-CA 90.83 ± 0.25 91.98 ± 0.18 90.72 ± 0.23 Zoo 86.92 ± 6.99 90.77 ± 8.85 90.77 ± 9.65 20newsW100 79.96 ± 0.77 81.04 ± 0.72 80.21 ± 0.75 Mushroom 99.78 ± 0.07 99.93 ± 0.03 99.61 ± 0.21 NTU2012 87.55 ± 1.52 89.07 ± 0.90 88.47 ± 1.92 ModelNet40 98.27 ± 0.21 98.18 ± 0.12 98.41 ± 0.11
D.3 Simplicial Higher-Order Datasets
Table 14 reports results introduced by Carrasco et al. (2025), a study conducted within the TopoBench framework, on real-world simplicial higher-order datasets. SCCNN achieves the highest overall performance, reaching $95.08\%$ accuracy on the NAME classification task, while also maintaining strong performance across other metrics. Simplicial complex-based networks (SCN, SCCNN, SaNN, GCCN) consistently outperform standard graph methods (GCN, GAT, GIN), with SCCNN exhibiting particularly stable results, indicated by a low standard deviation of $0.56$. These findings highlight the advantage of topological networks in modeling higher-order tasks, where conventional pairwise graph structures fall short in capturing complex relational patterns.
Table 14: Higher-order datasets. Results are shown as mean $\pm$ standard deviation. The best result is bold and shaded in grey, while those within one standard deviation are in blue-shaded boxes.
Model NAME ($\uparrow$) ORIENT ($\uparrow$) $\beta_{1}$ ($\uparrow$) $\beta_{2}$($\uparrow$) Graph GCN 42.14 $\pm$ 2.72 47.94 $\pm$ 0.00 46.86 $\pm$ 4.50 0.00 $\pm$ 0.00 GAT 18.09 $\pm$ 0.65 47.94 $\pm$ 0.00 7.45 $\pm$ 0.05 0.00 $\pm$ 0.00 GIN 76.14 $\pm$ 0.14 56.28 $\pm$ 0.45 88.13 $\pm$ 0.00 0.93 $\pm$ 1.21 Simplicial SCN 79.48 $\pm$ 1.36 69.55 $\pm$ 0.97 76.45 $\pm$ 3.06 5.45 $\pm$ 2.31 SCCNN 95.08 $\pm$ 0.56 86.29 $\pm$ 1.23 90.20 $\pm$ 0.20 65.82 $\pm$ 2.70 SaNN 81.76 $\pm$ 1.37 61.65 $\pm$ 0.55 88.46 $\pm$ 0.09 39.22 $\pm$ 2.80 GCCN 86.76 $\pm$ 1.27 76.60 $\pm$ 1.67 84.20 $\pm$ 4.80 41.82 $\pm$ 20.19 HOPSE-M 91.50 $\pm$ 1.45 80.68 $\pm$ 1.72 90.26 $\pm$ 0.55 71.69 $\pm$ 1.50 HOPSE-G 81.75 $\pm$ 1.26 62.17 $\pm$ 0.98 88.28 $\pm$ 0.08 35.37 $\pm$ 2.25
Appendix E Additional Dataset Details
To promote transparency, reproducibility, and ease of use, all dataset loading and preprocessing functionalities are encapsulated within the TopoBench library’s loader module (loader module). Most graph datasets are processed using the official torch_geometric loaders, which parse raw formats (e.g., .csv, .npz, or edge lists). The US-county-demos dataset is further adapted from the following repository (link). Higher-order hypergraph datasets are sourced from the repository of Chien et al. (2021) (link), while the MANTRA family datasets are adapted from (link) and integrated into TopoBench with consistent formatting adapted from torch_geometric, which stores the preprocessed datasets in the standardized PyTorch .pt files. This unified pipeline automates the full dataset preparation process and removes the need to access external repositories manually.
TopoBench includes all dataset licenses—where applicable—in the file located at the root of the repository and named third_party_licenses.txt. Additionally, a dedicated Datasets section in the README.md file provides references to the original source papers for each dataset included in the benchmark.
E.1 Graph datasets
Shared preprocessing: As emphasised in Appendix C.4–for demonstration purposes, only one fixed lifting is applied to transform graphs into each of the considered topological domains, leaving a domain-specific optimal lifting strategy beyond the scope of this paper. Specifically, a clique complex is used for simplicial lifting, cycle-based lifting is used for cellular domains, and $k$-hop lifting (with $k=1$) is used for hypergraphs. Feature projection is also applied, where the (n1)(n-1)-cell features are multiplied by the corresponding incidence matrices to generate $n$-cell features.
Cora, Citeseer, and Pubmed are adapted from the open-source Planetoid dataset collection available in the torch_geometric repository (link to dataset). Preprocessing: no additional preprocessing is applied beyond the shared one.
MUTAG, PROTEINS, REDDIT-BINARY, IMDB-BINARY, IMDB-MULTI, NCI1, and NCI109 are sourced from the open-source TUDataset collection (link to dataset). Preprocessing: no additional preprocessing is applied beyond the shared one.
ZINC is adapted from the open-source dataset available at (link to dataset). Preprocessing: node features are first transformed into one-hot encodings, after which the shared preprocessing is applied.
Amazon Ratings, Minesweeper, Roman Empire, and Tolokers are obtained using the HeterophilousGraphDataset loader from torch_geometric (link to dataset). Preprocessing: no additional preprocessing is applied beyond the shared one.
The US-county-demos dataset is taken from the official Cornell website (link to dataset). Preprocessing: the version used is already preprocessed as in Jia and Benson (2020), and no further preprocessing is applied beyond the shared one.
E.2 Higher-Order datasets
Higher-order hypergraph datasets are acquired from the repository of Chien et al. (2021) (link to dataset) and adapted to the benchmark pipeline to conform to the integrated torch_geometric format used in TopoBench. Additional information regarding the hypergraph datasets is provided and discussed in Appendix D.1.
The code for the MANTRA family datasets is adapted from the (https://github.com/aidos-lab/MANTRA) and integrated into TopoBench following the torch_geometric formatting.